If government tax policy requires Jasmin to pay $30,000 in taxes on annual income of $200,000 and Paul to pay $10,000 in tax on annual income of $100,000, then the tax policy is:

-Proportional

- optional

- regressive

-progressive
Answer #1

-progressive

(Average tax rate when income is 200,000 = Tax paid/Income = 30,000/200,000 = 0.15 = 15%
Average tax rate when income is 10,000 = Tax paid/Income = 10,000/100,000 = 0.10 = 10%
So, it is a progressive tax as average tax rate increases with increase in income.)